# SafeNav: Safe Path Navigation using Landmark Based Localization in a GPS-denied Environment

### v1

@@ V-D1 Average Percent Error in Trajectory Lengths
V-D2 Displacement Errors and Weighted Risk Scores
Table IV compares ADE, FDE, and AWRS across three path types $(P1$$P3)$. SafeNav-Centroid consistently outperforms SafeNav-CHull, with $1.67$ times lower ADE and $3.30$ times lower FDE, indicating higher accuracy and precision. Its lower final displacement errors suggest better alignment with the intended destination.
TABLE IV: Comparison of ADE/FDE, AWRS, and time for SafeNav-CHull and SafeNav-Centroid
@@ VI Conclusion and Future Works
In this paper, we present LanBLoc-BMM(EKF), an efficient navigation approach combining LanBLoc-based location measurement, BMM motion planning, and EKF integration. Building on this, we propose two safe path navigation algorithms, SafeNav-Chull and SafeNav-Centroid, which guide mobile troops along obstacle-free paths in GPS-denied battlefield environments. These algorithms combine geometric properties of the path segments (Convex Hull, Centroid) with a novel Risk-Aware RRT* and LanBLoc-BMM(EKF), achieving outstanding performance on localization and trajectory estimation.
Experimental results validate their effectiveness, showcasing their potential to significantly enhance the safety and operational efficacy of military forces in challenging and hostile environments. These approaches address the critical vulnerability of GPS reliance, ensuring troops can maneuver safely and efficiently in contested and hazardous areas. Future work will target reduced computational demand, enhanced adaptability to dynamic threats, and alternative localization methods for landmark-sparse settings.
Acknowledgement
This research was funded by the Army Research Laboratory under Cooperative Agreement Number W911NF2120261. The opinions in this paper are the authors’ and do not reflect the Army Research Laboratory’s or the U.S. Government’s official stance. The U.S. Government can legally reproduce and share this content for governmental use, despite any copyright claims.

### v2

@@ V-D1 Average Percent error in trajectory lengths of two approaches
V-D2 Displacement errors and Weighted Risk Scores of estimated trajectories
Table V compares ADE, FDE, and AWRS performances of two navigation algorithms across three different path types (P1, P2, P3). SafeNav-CHull shows higher errors suggesting less accurate trajectory estimation compared to the Centroid-based method. SafeNav-Centroid outperforms SafeNav-Chull, achieving the lowest ADE/FDE rates, indicating the highest accuracy and precision in trajectory estimation. The safeNav-Centroid algorithm is 1.67 times and 3.30 times better, respectively than the ADE and FDE of the Convex-Hull-based approach. The Centroid-based algorithms show better performance in terms of final displacement error than average displacement error which infers that these algorithms allow moving objects to reach precisely close to the intended safe destination.
TABLE V: Comparison of ADE/FDE, AWRS, and time for SafeNav-CHull and SafeNav-Centroid
@@ VI Conclusion and Future Works
In this paper, we first proposed an efficient navigation approach LanBLoc-BMM(EKF) using the LanBLoc as a location measurement model and BMM for motion planning, integrating EKF. Then, we proposed two efficient safe path navigation algorithms, SafeNav-Chull and SafeNav-Centroid, for guiding the mobile troops along the safe and obstacle-free path in GPS-denied battlefield environments. The proposed algorithms combine LanBLoc-BMM(EKF) and geometric properties of the path segments, such as Convex Hull and Centroid, integrated with a novel Risk-Aware RRT* algorithm. The proposed solutions have demonstrated outstanding accuracy in localization and trajectory estimation along defined safe paths. Experimentation and simulation results have validated the effectiveness of the proposed algorithms, showcasing their potential to significantly enhance the safety and operational efficacy of military forces in challenging and hostile environments. This approach addresses the critical vulnerability of GPS reliance, ensuring troops can maneuver safely and efficiently in contested and hazardous areas.
Our future research will focus on reducing computational demands, improving adaptability to dynamic obstacles and hazard scenarios, and exploring alternative localization strategies for environments lacking identifiable landmarks.